Bo and Sue make 20 drinks in 5 minutes. Sue can make 20 in 10 minutes. Joe can make 20 in fifteen minutes, if they work together how long will it take?
Answer by stanbon(75887)

Joe and Sue make 20 drinks in 5 minutes. Sue can make 20 in 10 minutes. Joe can make 20 in fifteen minutes, if they work together how long will it take?
----------------------------------
Sue DATA:
Time to make 20= 10 min ; Rate = (1/10) job/min
Joe DATA:
Time to make 20= 15 min ; Rate = (1/15) job/min
Together DATA:
Time to make 20= x min : Rate = (1/x) job/min
EQUATION:
rate + rate = rate together
1/10 + 1/15 = 1/x
25/150=1/x
25x=150
x=6 minutes (amount of time required working together)
Cheers,
Stan H.